if the sum three consicutive even numbers is equal to 258. find the number
Answer by mananth(16946)

let the numbers be x,x+2,x+4
...
x+x+2+x+4=258
3x+6=258
-6
3x=252
/3
x=252/3
x=84
....
84,86,88
adda them up too get 258